FANDOM


List of runsEdit

# Type Date Description
233 9/6 BGO6 with Na source, triggered on discriminated signal
240 9/6 BGO6 + NaI coincidence with Na Source
241 9/6 NaI with source, triggered on discriminated signal
242 9/6 BGO6 with Na source, without shaper
243 9/6 BGO1 without source, without target
286 9/8

Ch0 = NaI; Ch1 = BGO1;, Ch2 = BGO2; Ch3 = BGO5; Ch4 = BGO6;

Ch5 = PiStop; Trigger = (BGO1or2)and(BGO5or6)

288-298 9/8 useless
299 9/8 Same as 286, but Ch5 = (BGO1or2)and(NaI)
300 9/8

Ch0 = NaI; Ch1 = BGO1;, Ch2 = BGO2; Ch3 = BGO5; Ch4 = BGO6;

Ch5 = nothing; Ch6 = PiStop; Trigger = (BGO1or2)and(NaI)

301 9/9 Same as 300, but Ch5 = S2andS3andS5andS6and(notS4);
302-307 9/8 some tests, don't use
308 9/9

Some DAQ data.

Ch0 = NaI; Ch1 = BGO3, Ch2 = BGO2; Ch3 = BGO5; Ch4 = BGO6; Ch6 = PiStop w=150ns; Trigger (BGO2or3)and(BGO5or6orNaI) w=2/mus

Not sure about recorded data.

309 9/9 Data taking, same DAQ channels as 308, 155 degrees as it was initially.
310 9/9 Same as 309, 175 degrees.
312 9/9 Same as 310. No target.
313 9/9 No beam, check rates.
314 9/9 Test run, config. same 312
315 9/9 Test run to check height of BGO signal, not for offline analysis
316 9/9

Test run to look at spectre of PiStop, not for offline.

317 9/9

Test with attenuated calo signals to avoid saturation in the logic box. Trigger on 2 calo coinc. Search for pi->yy signal and trigger-calo-delay.

318 9/9 Same as before
319 9/9 Same same, but different
320 9/9

Energy calibration for the calorimeters, that were put into the beam line.

No absorber, no target, at 0.5x170MeV

Ch0 = NaI; Ch1 = BGO8, Ch2 = BGO7; Ch3 = BGO5; Ch4 = BGO6; Ch6 = PiStop w=150ns; Trigger (BGO2or3)OR(BGO5or6orNaI) w=2/mus

321 9/9 same but: 0.8x170MeV
322

at 170 MeV

323 9/9 same as before
324 9/9 Beam on BGOs 7 & 8, beam energy 1.3*170MeV=221MeV
325,326 9/9 Let's pretend these didn't happen, eh?
327 9/9 Beam on BGOs 7 & 8, beam energy 85MeV (05.x before)
328 9/9 Moved the BGOs to beam on 5&6. Same energy as last run. Peak at 186900+-976.5
329 9/9 Changed energy to 170MeV. Should get a pretty good peak around 305500+-350.8.
330 9/9 Changed energy to 136MeV. Peak around 258400+-699.8.
331,332,333 9/9 Discard.
334 9/9 Changed energy to 119MeV. Peak around 249k. But barely because the beam stopped. The data file may be corrupted...
335 9/9 Re-arranged stuff. Movable ecal at 155deg closest to beam. Energy=1.3*170MeV=221MeV. Didn't work though =( Perhaps something was wrong with the co-incidence; nevertheless data are useless.
336 9/9 Same again, not really useful.
337 9/9-10/9 Same again, look at channel 3, BGO 5, and channel 4, BGO6.
338 9/10 ??
339 9/10

221MeV, angle changed to 158 degrees (beam side)

Channels: NaI,BGO3,BGO2,BGO5,BGO6,not in use,PI-STOP

The 2 different BGOs on each side lead to 3 different angles covered per run if used in coincidence (offline analysis). See logbook for more details.

340-358 9/10 daq test, do not use
359 9/10 same as 339
364 9/10 160deg
368 9/10

156deg, higher BGO threshold. from 314 to 1517 units (to go to mV -> divide by 10)

lower total rate, but about the same for high energies => still to check!

369 9/10 daq test, do not use
370 9/10 170deg. Expect to see significantly lower number of  high energy events. For angular dependency of the rate see logbook.
371-374 9/10 test runs with small statistics
375 9/10 Same setup as run 370, time: 8:18-9:26
376 9/10 Same setup as run 370, time: 9:26-10:18
377 9/10

Changed Setup: moved BGOs back to 156.5deg. Plan to aquire high statistics because offline analysis doesn't show any signals yet. 156.5deg yields highes yy-rate. time: 10:47-11:50

378 9/10 garbage
379 9/10 same as 377, time 11:50 - 13:13
380 9/10 same as 377, time 13.15-14:22
381 9/10 same as 377, time 14:23-15:10
382 9/10 same as 377, time 15:10-15:15
383 9/10

garbage (beam shutter was closed)

384 9/10

same setup as 377 but with TARGET REMOVED, time 15:27-16:18

385 9/10 see 384, time 16:19-18.41
386 9/10 see 384, time 18.41-??
400/401 11/9 Roughly 1 hr combined, after changing RF delay to 5ns.
402 11/9 Discard.
403 .
404 . moved s4 to 11.5cm from target. removed target.
405 . With target, vertical position changed. Started at 0247. Ended at 0325(?) (its in the log book)
406 . Moved slits to 450. Then decided this was sub-optimal. Started at 0429. Stopped at 0440.
407 . Moved slits back to 500. Run started at 0443. Stopped temporarily at 0552.
408 . Minor problem with labview, no actual data recorded.
409 . Continuation of 407, at 0552
410 . Ditto, 0632. 0658
411 9/11 Ditto, 0707. 0824
412 9/11 Clock signal onto BGO2356NaI cables
413 9/11 continuation of 407, 9:1210:46
414-424 9/11 garbage
425

NEW SETUP: Trigger on Pi-stop - calo - coinc. , time 16:27-17:34

426 short run, garbage
427
CH2
9/11 NEW SETUP: Removed S4 and S1, Trigger: incoming Pi in coinc with calo, time 18:06 - 19:00
428
CH2
9/11 Same as 427, time 19:00 - 20:07
429
CH2
9/11 Same as 427, time 20:14 - 21:14
430
CH2
9/11 Same as 427, time 21:14 - 22:12
431
CH2
9/11

Same as 427, time 22:12 - 23:12

432 9/11 garbage (DAQ malfunction)
433
CH2
9/11 same as 427, time 23:18
434
C
9/12 Carbon Target measurements
435
C
9/12 Carbon Target measurements
436 ? 9/12 Carbon Target - Computer restarted, maybe dmg files
437
C
9/12 Carbon Target measurement
438
C
9/12 Carbon Target measurement
439
NT
9/12 No target,  Time started: 0544.
440 9/12 measurement 440 garbage (DAQ crash after 1 min data taking).
441
NT
9/12
442
NT
9/12 Time finished: 0900.
443 12/9 CH2 put back in and started at: 0910. DISCARD Stopped at 0917 due to dodgy activity with the power supply. signals jumping around all over the place.
444-446 Discard.
447
CH2
Start CH2 at 0926. Pause at 1000.
448-450 Discard.
451
CH2
Continue at 1002 to ~1100
452
CH2
continue at ~1100 to 1200
453
CH2
continue at 1200 to 1244
454
CH2
continue at 1245 to 1356
455
CH2

continue at 1356 to 1511

456
CH2

continue at 1511 to 1612

457
CH2

continue at 1612 to 1712

458
CH2

continue at 1712 to 1815

459
CH2

continue at 1815 to 1915

460
CH2

continue at 1915 to 2141. Broken run, do not use.

461
CH2

continue at 2141 to 2215. Broken run, do not use.

462
CH2

continue at 2215 to 2315

463 CH2 continue 23.15 to 24.00
464 C 00.10 - 00.55ish
465 no beam :(
466-468 C 11.55-13:54
469 Trash Last Run, stopped by Myriam ;)


Analysis of the histograms:

Run # 511 kev σ [%] res peak hight 1.2 mev σ [%] res peak hight peak pos ratio peak heigt ratio
233 11.54 1.77 15.3 162 27.36 2.84 10.4 77.8 2.37 2.1
240 11.14 0.45

4

241 10.98 0.46 4.2 111 26.53 0.99 3.7 36.3 2.41 3.1
242 7.94 1.15 14.5 662 20.16 1.94 9.6 127 2.53 5.2

Ad blocker interference detected!


Wikia is a free-to-use site that makes money from advertising. We have a modified experience for viewers using ad blockers

Wikia is not accessible if you’ve made further modifications. Remove the custom ad blocker rule(s) and the page will load as expected.